# Inverse design of potential metastructures inspired from Indian medieval architectural elements

### v2

@@ Results
In figure 5, we have selected a primary honeycomb lattice structure interlaced with a single honeycomb unit (LS1) at the centre. Our objective is to study the effect of mass insertion in the central honeycombs on the bandgap characteristics of the metastructure. The metastructure is chosen as per the ease of mass insertion and enables us to observe changes in the attenuation band, which are influenced by local resonances occurring at frequencies that vary with the insertion location. Metallic hexagonal inserts are made up of structural steel (as shown in Supplementary Fig. S5), and their detailed properties and composition are shown in Supplementary table S3. Given that the lattices are fabricated through 3D printing using Polylactic Acid (PLA), and subsequently metallic inserts of much heavier density are used inside the lattice, these inserts function as local resonators due to the density difference between PLA and steel. These characteristics make the metastructure well-suited for vibration control applications. The presence of inserts significantly influences the wave transmission patterns. Furthermore, the strategic placement of the inserts are considered in such a manner that the inserts themselves make a pattern inside the lattice structure. By changing their location, we can control the bandgap for planar excitation in 1D. In Figure 5(a), we present the effect of random vibration through the simple lattice, while in Fig. 5(b) we explore the effect of wave transmission in broken symmetry of diamond pattern where the middle row of the lattice is fully filled while the top and bottom row is filled with metallic inserts with a definite pattern. In Fig. 5(c), the effect of wave transmission in the cross-shape pattern is shown where the inserts are arranged alternatively in the middle row of lattice and the top and bottom row is completely filled. In Fig. 5(d) lattices are filled in such a manner that it forms an arrow indicating towards the right direction. Figure 5(e), represents the diamond pattern where the middle row is fully filled; however, the top and bottom rows are alternatively filled. In Fig. 5(f), the top and bottom rows are fully filled and the middle row is fully empty. In Fig. 5(g), the top and bottom rows are fully filled; however, the middle row is filled such that there is a spacing of two units between the consecutive inserts. In Fig. 5(h), the top row is alternatively filled starting with a blank cell while the bottom row is also alternatively filled starting with a filled cell, and the middle row is fully filled except for the first and the last cell representing upward pointing arrow. In Fig. 5(i), the top and bottom rows are alternatively filled while the middle row is filled with two-unit spacing. Through the study of dynamic response corresponding to harmonic excitation for all these combinations, we aim to find out which pattern location is more sensitive and affects more at a specific frequency range of wave transmission. We observed that generally, the middle row exhibits higher sensitivity compared to the top and bottom rows. Additionally, by arranging different combinations of inserts, we achieved up to 48 % bandgap enlargement, multiple bandgaps, and the ability to shift the bandgap across the desired frequency range. Further, the experiments corresponding to each iteration are carried out to validate our simulation results by using a Laser Doppler Vibrometer (LDV). Random vibration up to 10 KHz is performed using LDS shaker with the gain set as unity. Detailed experimental setup can be seen in Supplementary Fig. S10. As may be observed in Fig. 5, our experimental results show good agreement with the simulation. The printing pattern, and infill density (presented in Supplementary Table S2), also affect the attenuation band.
Bandgap characteristics of honeycomb based interlaced structures
Figure 6: (a) A comparison of bandgap at different frequency levels of different structures interlaced in the honeycomb structure. (b), (d) The comparison of static performance is obtained through analytical computations of different structures. The displacement control mechanism is set as 5 mm compression of initial height of samples (c) Comparison of bandgap in different order (n=0,1,2) of the fractal structure.
Comparative analysis of the effect of different interlaced structures on bandgap shift (Fig. 6(a)) is carried out in various frequency ranges. The band gap occurs at a frequency band comparatively higher than that of the simple honeycomb structure. Notably, for the LS3 structure bifurcation in bandgap occurs, and for the last structure (LS9) multiple bandgaps occur both for the higher and lower frequency ranges. Furthermore, we have compared the bandgap range of various interlaced honeycomb unit cells (Fig. 6b) along with the force-deflection relationships. The force-deflection relationships are linear for all the patterns except for LS9.
Figure 6(c) illustrates the fractal structure of different order (n=0,1,2) and their corresponding bandgap shifts. In different structures, as the fractal order increases, the bandgap shifts to higher frequency ranges, and for further increase in fractal order multiple bandgaps occur. Furthermore, on the right side we have represented the force-deflection relationship for the fractal structure (see Fig. 6d). With the increase in the fractal order, the structure becomes stiffer, with 196.32% increase from zeroth to first-order fractal and 33.61% increase from first to second-order fractal. From this study, we can conclude that with increasing the structural complexity inside the structure, the bandgap starts to shift towards the higher frequency range, and for higher-order fractals, it also bifurcates which is verified by taking different fractal structures inside the honeycomb and explained in the Supplementary Fig. S8.
AI-driven Inverse Design of Metastructures
@@
With the phenomenal expansion potential of the geometry of metastructures, it is possible to use AI efficiently for both forward analysis in predicting the bandgaps as well as inverse design where for a given bandgap a suitable pattern may be obtained. The complete flow diagram of a possible AI-based inverse design of metastructures is depicted in Fig. 7. The entire method comprises two components: (i) the forward analysis model and (ii) the inverse design model. The forward analysis model acts as a surrogate model for training the inverse design model. It is trained independently for the forward analysis of the metastructures with known transmissibility.
@@
Forward Analysis Model
The forward analysis model consists of repeated blocks of geometric feature extraction modules capable of learning complex metastructure geometry having primary, secondary, and tertiary complexity. Each geometrical feature extraction module consists of a multi-scale geometric attention mechanism having multiple heads of FEM-inspired spatial attention (FSA) blocks. The FSA is designed to learn the dependencies of adjacent connections at different scales using different kernel sizes of convolutional modules. Each FSA consists of an input convolutional layer for feature transformation followed by a spatial attention block [61, 62]. The complete block diagram of the forward analysis block, multi-scale GA module, and the FSA block are provided in the supplementary Figures S13 and S14. The forward analysis model was implemented with 12 layers of geometric feature extraction blocks and 8 heads of FSA in each block. The kernel sizes of the input convolutional layer in 8 FSA heads are selected as $(1,1,1,3,3,3,5,5)$. This combination of kernel sizes puts more attention on the smallest structures (tertiary), than the secondary, and the primary structures. The output of the geometric features extraction module is processed by a regression head that transforms the geometric features to the transmissibility of dimension $(1000,1)$. The detailed block diagram and the descriptions are provided in the supplementary file: Fig. S15 and Section E.1.
Inverse Design Model
The inverse design model is developed to obtain the metastructure geometry using transmissibility plots for the desired bandgap. The detailed block diagram of the inverse design model is provided with the supplementary materials (Figures S16 and S17). The key contributions of the inverse design model are the conversion of transmissibility into latent space using a multi-head 1-D self-attention mechanism, followed by constructing the interlaced patterns using a multiscale residual network. The 1-D self-attention mechanism captures the sequential dependencies of the transmissibility spectrum, [63] and generates a rectified latent space of shape $(3\times 128\times 256)$. The latent space is further extended to a new higher space using reflection padding and then non-linearly down-sampled back to the same dimension using multiple consecutive convolutional layers with ReLU activation.
@@
Now the transformed latent space is passed through a multiscale residual network having three parallel convolutional paths with three different kernel sizes: $(1\times 1)$, $(3\times 3)$, and $(5\times 5)$ as shown in Fig. S17(b). These sizes have been selected based on the institution to transform the latent space into the three levels of interlaced structural geometry. The smallest kernel is responsible for learning the tertiary structures, the medium size helps to learn the secondary structural geometry. Similarly, the largest value $(5\times 5)$ is responsible for learning the primary geometry. The final output of the multi-scale residual network is the interlaced patterns. It is now upsampled using three layers of 2-D transposed convolution and then transformed non-linearly using 2-D convolution with hyperbolic tangent (Tanh) activation function to produce the final geometry.
Training
For the training of the proposed AI-based inverse design of the metastructures, a dataset having 720 samples of different metastructures is prepared using different interlaced structures as shown in supplementary Fig. S12. In the unit cell of these structures, the primary structure is a honeycomb and the secondary structures are different combinations of honeycomb, star, circle, hierarchical honeycomb, and auxetic interlaced structure. Then the corresponding transmissibility is calculated using FEM analysis.
@@
The samples are divided into two portions: training and validation with a ratio of 90:10. The AI-based inverse design mechanism is trained in two phases.
@@
    Forward Analysis Model: In the first phase, the forward analysis model is trained using the training samples to predict the transmissibility of a given metastructures. The training loop is depicted in Fig. 7(a). The comparison of the predicted transmissibility obtained by the forward analysis model and the true transmissibility obtained by the FEM analysis are compared in Fig. 8.
@@
    Inverse Design Model: Once the forward analysis model is fully trained, it is used as a surrogate model for the training of the inverse design model as displayed in Fig. 7(b). The loss function $(\mathcal{L})$ for the weight update of the inverse design model is constructed as provided in equation (1)
@@
    where, $\beta_{1}$ and $\beta_{2}$ are the penalty terms for the inclusion of losses $\mathcal{L}_{2}$ and $\mathcal{L}_{3}$. The values of $\beta_{1}$ and $\beta_{2}$ were selected as 0.1 and 0.6, respectively.
@@
    $\mathcal{L}_{1}\;=\;\text{MSE}$: mean square error between predicted transmissibility on generated metastructure and the transmissibility obtained from the FEM method on original metastructure.
    $\mathcal{L}_{2}\;=\;1-\text{SSIM}$: structural similarity index measure between the generated metastructure and the original metastructure [64].
    $\mathcal{L}_{3}\;=\;\text{MMD}$: maximum mean discrepancy between the generated metastructure and the original metastructure (MMD) [65].
    Further details of these losses are provided in the supplementary material, Section E.3.
Validation of AI-model for metastructure design
The metastructure produced by the inverse design model contains colored noises due to training with a small value of batch size (small batch size was chosen due to the limited number of samples in the dataset). Therefore, a noise removal filter is applied to produce a clear image of the obtained metastructure during validation. The algorithm for the noise removal is provided in supplementary material, Section F. The AI-based inverse design of metastructures is validated under two cases:
@@
Figure 8: Validation results for AI-based inverse design of metastructure. (a) Original metastructure of the test sample. (b) Input transmissibility obtained by FEM for the original metastructure. (c) Metastructure predicted (generated) by the inverse design model using the input transmissibility. SSIM = 0.991, closer to unity validates the exact match with the original image. (d) Comparing the transmissibilities of the original metastructure obtained from the forward model and FEM. (e) Comparing the transmissibilities of the generated metastructure obtained from the forward model and FEM.
Validation on known metastructure as test sample:

### v3

@@ 3.1 Performance measurement of lattice with resonators
In figure 7, we have selected a primary honeycomb lattice structure interlaced with a single honeycomb unit (LS1) at the centre. Our objective is to study the effect of mass insertion in the central honeycombs on the bandgap characteristics of the metastructure. The metastructure is chosen as per the ease of mass insertion and enables us to observe changes in the attenuation band, which are influenced by local resonances occurring at frequencies that vary with the insertion location. Metallic hexagonal inserts are made up of structural steel (as shown in Supplementary Fig. S5), and their detailed properties and composition are shown in Supplementary Table S3. Given that the lattices are fabricated through 3D printing using Polylactic Acid (PLA), and subsequently metallic inserts of much heavier density are used inside the lattice, these inserts function as local resonators due to the density difference between PLA and steel. These characteristics make the metastructure well-suited for vibration control applications. The presence of inserts significantly influences the wave transmission patterns. Furthermore, the strategic placement of the inserts is considered in such a manner that the inserts themselves make a pattern inside the lattice structure. By changing their location, we can control the bandgap for planar excitation in 1D. In Figure 7(a), we present the effect of random vibration through the simple lattice, while in Fig. 7(b) we explore the effect of wave transmission in broken symmetry of diamond pattern where the middle row of the lattice is fully filled while the top and bottom row is filled with metallic inserts with a definite pattern. In Fig. 7(c), the effect of wave transmission in the cross-shape pattern is shown where the inserts are arranged alternatively in the middle row of the lattice and the top and bottom rows are completely filled. In Fig. 7(d) lattices are filled in such a manner that it forms an arrow indicating the right direction. Figure 7(e), represents the diamond pattern where the middle row is fully filled; however, the top and bottom rows are alternatively filled. In Fig. 7(f), the top and bottom rows are fully filled and the middle row is fully empty. In Fig. 7(g), the top and bottom rows are fully filled; however, the middle row is filled such that there is a spacing of two units between the consecutive inserts. In Fig. 7(h), the top row is alternatively filled starting with a blank cell while the bottom row is also alternatively filled starting with a filled cell, and the middle row is fully filled except for the first and the last cell representing upward pointing arrow. In Fig. 7(i), the top and bottom rows are alternatively filled while the middle row is filled with two-unit spacing. Through the study of dynamic response corresponding to harmonic excitation for all these combinations, we aim to find out which pattern location is more sensitive and affects more at a specific frequency range of wave transmission. We observed that generally, the middle row exhibits higher sensitivity compared to the top and bottom rows. Additionally, by arranging different combinations of inserts, we achieved up to 48% bandgap enlargement, multiple bandgaps, and the ability to shift the bandgap across the desired frequency range. Further, the experiments corresponding to each iteration are carried out to validate our simulation results by using a Laser Doppler Vibrometer (LDV). Random vibration up to 10 KHz is performed using LDS shaker with gain set as unity. Detailed experimental setup can be seen in Fig. 6. As observed in Fig. 7, our experimental results show good agreement with the simulation. The printing pattern, and infill density (presented in Supplementary Table S2), also affect the attenuation band.
4 AI-driven Inverse Design of Metastructures
@@
With the phenomenal expansion potential of the geometry of metastructures, it is possible to use AI efficiently for both forward analysis in predicting the bandgaps as well as inverse design where for a given bandgap a suitable pattern may be obtained. The complete flow diagram of the proposed AI-based inverse design of metastructures is depicted in Fig. 8. The entire method comprises two components: (i) the FEM-inspired spatial attention (FSA)-based forward analysis model and (ii) the multiscale Gaussian self-attention (MGSA)-based inverse design model. The forward analysis model acts as a surrogate model for training the inverse design model in the tandem structures of neural networks. The forward analysis model is trained independently using the training samples of metastructure and its transmissibility for the prediction of transmissibility for new metastructures.
@@
Figure 9: Detailed block diagram of the FSA-based forward analysis model: (a) Complete block diagram of the forward model showing the transformation of input metastructure to output transmissibility. The orange box shows the geometric attention (GA) model responsible for capturing the geometrical dependency of vibration mitigation. Grey boxes show the 8 heads of the FEM-inspired spatial attention (FSA) module. (b) Detailed diagram of FSA module. The kernel size of the input convolutional block ($f^{\alpha\times\alpha}$) is different for each FSA head for capturing the inter-dependencies of adjacent connections in the metastructure geometry.
4.1 Forward Analysis Model
The forward analysis model consists of $N$ blocks of geometric feature extractor (GFE), followed by a regression head to process the output of GFE and the material properties to predict the transmissibility of the metastructure. The GFE module is designed on the concept of finite element model (FEM) with the capability of learning the neighborhood dependencies in the complex metastructure geometry having primary, secondary, and tertiary complexity. Each GFE module consists of a multi-scale geometric attention (GA) mechanism having multiple heads of FEM-inspired spatial attention (FSA) blocks. Layer normalization (LN) before the GA and after the summing block is provided to stabilize the training process by reducing internal covariate shifts. Dropout after the GA block is included for the weight regularization. The complete forward model, along with the detailed diagram of FSA is shown in Fig. 9.
Each FSA consists of an input convolutional layer with a different kernel size followed by a spatial feature attention with learnable weights. The kernel sizes of the input convolutional layer in different FSA heads are selected as $\alpha=1,3,$ and $5$ depending on the complexity of the input geometry. If $x_{f}=f^{\alpha\times\alpha}(x)$ is the output of the input convolution, the geometrical feature output based on neighborhood dependencies is obtained using the following equation (1).
$x_{a}^{s}=f^{\beta\times\beta}\left(\left[w_{1}F^{s}_{avg}(x_{f}),\;\;w_{2}F^{s}_{max}(x_{f})\right]\right)\implies x_{o}=x_{f}\otimes\sigma\left(x_{a}^{s}\right)$ (1)
where, $f^{\beta\times\beta}(.)$ denotes the convolution with kernel size $\beta$, $\sigma(.)$ denotes the sigmoid operation, and $\otimes$ represents the matrix multiplication. $w_{1}$ and $w_{2}$ are learnable weight parameters, initialized with unity. The kernel size $\beta$ is a constant value equal to 7 for our case. Learnable weights in the FSA block are optimized to learn the inter-dependencies of adjacent connections in the spatial dimension. Overall, the whole process attempts to mimic the finite element model to map the vibration spectrum with the metastructure geometry.
[…]
The output of the GFE module is processed by a regression head that transforms the geometric features to the transmissibility of dimension $(1000,1)$. The fully connected (FC) layers in the regression head are associated with Parametric Rectified Linear Units (PReLU) activation. The PReLU activation function facilitates learnable non-linearity to map the highly non-linear nature of the transmissibility. The second FC layer is provided with material properties along with geometric features. These material properties are modulus of elasticity $(E)$, Density $(\rho)$, Poisson’s ratio $(\nu)$, and material damping $(d)$. The output of the last-second FC layer is applied to layer normalization (LN) to stabilize the learning. Finally, the last FC layer with linear activation converts the features to the transmissibility having dimension (1000, 1)10001(1000,\;1)( 1000 , 1 ).
Figure 10: Detailed block diagram of the inverse design model. (a) Complete block diagram of inverse design model transforming input transmissibility to metastructure geometry. (b) Detailed block diagram of spectrum self-attention responsible for capturing the sequential dependencies of the transmissibility and generating a rectified latent space. The spectrum position encoding is specially designed using the Gaussian function to emphasize the specific frequency range in the non-linear spectrum. (c) Detailed block diagram of the multiscale residual network. The kernel sizes of the convolutional blocks in three channels are $(1\times 1)$, $(3\times 3)$, and $(5\times 5)$, selected to transform the latent space into the three levels of interlaced structural geometry.
4.2 Inverse Design Model
The inverse design model is developed to obtain the metastructure geometry using transmissibility data for the desired bandgap. The key constituents of the inverse design model are Gaussian self-attention to convert transmissibility to latent space, down sampling network, multiscale residual network to produce interlaced complex structure, and finally upsampling through transpose 2D convolution as shown in Fig 10(a). The Gaussian self-attention is a multi-head self-attention mechanism with a Gaussian function for 1D spectrum position encoding. It captures the sequential dependencies of the transmissibility and generates a rectified latent space of shape $(3\times 128\times 256)$, shown in Fig 10(b). Since the transmissibility is the amplitude versus frequency spectrum, a position encoding must be added to represent the absolute position of the amplitude against the log-scale frequency bin. Therefore, a learnable position encoding is defined using sine and cosine of Gaussian functions of log-scale frequency bin to represent the even and odd positions, shown in Equations (2) and (3), respectively.
$\text{PE}(2p)$ $=$ $\sin\left(\exp\left({-\frac{(\log(p)-\mu_{0})^{2}}{2\sigma_{0}^{2}}}\right)\right)$ (2)
$\text{PE}(2p+1)$ $=$ $\cos\left(\exp\left({-\frac{(\log(p)-\mu_{0})^{2}}{2\sigma_{0}^{2}}}\right)\right)$ (3)
where $p$ = position index, $\mu_{0}$ and $\sigma_{0}$ are the learnable parameters for the dynamic allocation of the mean and standard deviations of the Gaussian distribution for 1D position encoding. $\mu_{0}$ is initialized with half of the sequence length, whereas $\sigma_{0}$ is initialized with unity to ensure small spread. The optimization of these parameters enables the Gaussian function to emphasize the specific frequency range in the non-linear spectrum.
@@ 4.2 Inverse Design Model
The latent space generated by the spectrum self-attention is further extended to a new higher space using reflection padding with a reflection of 3 pixels and then non-linearly down-sampled back to the same dimension using multiple consecutive convolutional layers with ReLU activation. Now the transformed latent space is passed through a multiscale residual network having three parallel convolutional paths with three different kernel sizes: $(1\times 1)$, $(3\times 3)$, and $(5\times 5)$, shown in Fig. 10(c). These sizes have been selected based on the institution to transform the latent space into the three levels of interlaced structural geometry. The smallest kernel is responsible for learning the tertiary structures, the medium size helps to learn the secondary structural geometry. Similarly, the largest value $(5\times 5)$ is responsible for learning the primary geometry. The final output of the multi-scale residual network is the interlaced patterns. It is now up-sampled using three layers of 2D transposed convolution and then transformed non-linearly using 2D convolution with hyperbolic tangent (Tanh) activation to produce the final geometry.
4.3 Dataset preparation
Figure 11: Metastructure of primary geometry as honeycomb and secondary geometry (interlaced structures) including honeycomb, auxetic, star, circle, and hierarchical honeycomb used to design different combinations of metastructures for data set generation using numerical simulation.
A transmissibility dataset was generated for various unique metastructures, having multiscale geometrical designs. These metastructures featured a primary honeycomb structure with different secondary interlaced structures: honeycomb, star, circle, hierarchical honeycomb, and auxetic as shown in Fig.11. By integrating these secondary structures with the primary honeycomb, a variety of 720 distinct metastructures having different transmissibilities were created. The transmissibility of each structure was calculated by post-processing the frequency response obtained from Finite Element (FE) simulations conducted in Ansys. Therefore, the training dataset contains 720 samples of metastructure and corresponding transmissibility.
4.4 Training
@@
The proposed frameworks of forward analysis and the inverse design models were trained on the metastructure-vs-transmissibility dataset as described above. The total samples in the dataset were divided into two portions: training and validation with a ratio of 90:10, resulting in 648 samples in the training set and 72 samples in the validation set. The training of the proposed framework was accomplished in the following two phases.
@@
    Forward Analysis Model: In the first phase, the forward analysis model was trained using the training samples to predict the transmissibility of a given metastructures. The training loop has been depicted in Fig. 8(a). The weight optimization was achieved by minimization of the mean square error between the predicted transmissibility and the true transmissibility. The model was trained for 300 epochs using the Nesterov-accelerated Adaptive Moment Estimation (NAdam) as the optimization algorithm with dynamic allocation of learning rate by the Cosine Annealing learning rate scheduler (details can be found in [63] and [64]). The comparison of the predicted transmissibility obtained by the forward analysis model and the true transmissibility obtained by the FEM analysis are compared in Fig. 12.
@@
    Inverse Design Model: The trained forward model, obtained from the first phase of the training was used as a surrogate model for the training of the inverse design model as displayed in Fig. 8(b). The weight optimization was accomplished again using the NAdam optimization algorithm and dynamic allocation of learning rate using the Cosine Annealing learning rate scheduler (details can be found in [63] and [64]). A composite loss function $(\mathcal{L})$ for the weight update of the inverse design model is defined in Eq. (4)
@@
    where, $\mathcal{L}_{1}$ = mean square error (MSE), $\mathcal{L}_{2}$ = structural similarity loss (SSL), and $\mathcal{L}_{3}$ = maximum mean discrepancy (MMD). $\beta_{1}$ and $\beta_{2}$ are the penalty terms for the inclusion of losses $\mathcal{L}_{2}$ and $\mathcal{L}_{3}$. The values of $\beta_{1}$ and $\beta_{2}$ were selected as 0.1 and 0.6, respectively. The loss terms MSE, SSL, and MMD are defined as follows:
@@
    1. Mean square error (MSE): MSE represents the mean deviation of the predicted transmissibility of the generated metastructure and the true transmissibility of the original metastructure. If $a_{i}\in\mathcal{A}$ and $\hat{a}_{i}\in\hat{\mathbf{A}}$ are $i^{th}$ data point in the true and predicted transmissibilities, the MSE is defined in Eq. (5).
      $\mathcal{L}_{1}(\mathcal{A},\hat{\mathcal{A}})=\frac{1}{n}\sum_{i=1}^{n}\left(a_{i}-\hat{a_{i}}\right)^{2}$ (5)
    2. Structural similarity loss (SSL): SSL represents the structural misalignment of the generated metastructure and the original metastructure. It is defined in terms of structural similarity measure as shown in Eq. (6)
      $\mathcal{L}_{2}(\textbf{X},\hat{\textbf{X}})=1-\frac{\sigma_{\textbf{X}\hat{\textbf{X}}}}{\sigma_{\textbf{X}}\sigma_{\hat{\textbf{X}}}+c_{0}}$ (6)
      where $\sigma_{\textbf{X}\hat{\textbf{X}}}$ = covariance between input image X and the generated image $\hat{\textbf{X}}$, $\sigma_{\textbf{X}}$ = local variance (standard deviation) of the input image X, $\sigma_{\hat{\textbf{X}}}$ = local variance (standard deviation) of the generated image $\hat{\textbf{X}}$, and $c_{0}$ = small constant for division stabilization. The inclusion of SSL with a penalty factor of $\beta_{1}$ helps the algorithm to force the weight update in the direction to attain the structural similarity between the generated metastructure and the original metastructure. Since the SSL is the opposite of the structural similarity, therefore, $(1-SSL)$ is used to measure the similarity score of the generated image to the original image [65].
    3. Maximum mean discrepancy (MMD): MMD provides the statistical difference (non-parametric distance) between two probability distributions on reproducing kernel Hilbert space (RKHS) [66, 67]. Therefore, MMD can be used for comparing the statistical distributions of the two images by mapping them into a higher-dimensional feature space using a kernel function. If $x_{i}\in\textbf{X}$ and $\hat{x}_{j}\in\hat{\textbf{X}}$ are $i^{th}$ and $j^{th}$ data points in the original and the generated image space, respectively, then the MMD of the two distribution in RKHS is defined in Eq. (7)
      $\mathcal{L}_{3}(\textbf{X},\hat{\textbf{X}})=\left\|{\frac{1}{{n}}\sum\limits_{i=1}^{{n}}{f\left({{x_{i}}}\right)}-\frac{1}{{\hat{n}}}\sum\limits_{j=1}^{{\hat{n}}}{f\left({{\hat{x}_{j}}}\right)}}\right\|_{\mathcal{H}}$ (7)
      where, $f(.)$ denotes the kernel function $f:X,X\to{\mathcal{H}}$, ${\mathcal{H}}$ be the universal RKHS, $n$ and $\hat{n}$ be the number of pixel points in the original and the generated image X and $\hat{\textbf{X}}$ respectively. The MMD, calculated between the original metastructure and predicted metastructure, with a penalty factor of $\beta_{2}$, forces the training algorithm to minimize the discrepancies of the generated image compared to the original image.
    The inclusion of SSL and MMD in the loss function helps the training algorithm to optimize the inverse design model to learn the original geometry and makes it less dependent on the performance of the forward analysis model in the tandem network.
@@
Figure 12: Validation results for AI-based inverse design of metastructure. (a) Original metastructure of the test sample. (b) Input transmissibility obtained by FEM for the original metastructure. (c) Metastructure predicted (generated) by the inverse design model using the input transmissibility. The similarity score = 0.991, closer to unity validates the exact match with the original image. (d) Comparing the transmissibilities of the original metastructure obtained from the forward model and FEM. (e) Comparing the transmissibilities of the generated metastructure obtained from the inverse model and FEM.
4.5 Validation of AI-model for metastructure design
The metastructure produced by the inverse design model contains colored noises due to training with a small value of batch size (small batch size was chosen due to the limited number of samples in the dataset). Therefore, a noise removal filter is applied to produce a clear image of the obtained metastructure during validation. The algorithm for the noise removal is provided in supplementary material, Section E. The AI-based inverse design of metastructures is validated under two cases:
4.5.1 Validation on test samples of metastructure: